Please help!! Need answers ASAP with clear and concise explanations

1. 68Ga has a half-life of 68 minutes. If the activity that is to be administered to the patients is 15 mCi and the patient is to be injected at 10:00am, and if the radioisotope is synthesized at 7:15 am, what activity of 68Ga must there be at the time of synthesis before being shipped to the hospital?

2. Calculate the resolution of a SPECT system given the following characteristics. The intrinsic resolution of the crystal, which is 1 cm thick, and electronics is 4 mm. The collimator has the dimensions in which the length of the holes that are between septa is 5 mm (i.e., the thickness of the collimator), the thickness of each septa is 1 mm and the distance between individual septa is 1 mm. The distance of the source from the collimator is 1 cm. Assume that the collimator material attenuates 100% of the photon that interact with it.

3. A certain stable neutral element has 29 electrons. What is the element? A radioactive isotope of this element has 6 more neutrons than electrons. What is the isotope? This radioactive isotope decays and the resulting isotope has one fewer proton. What is the new element?
Answer #1

1.

T = half life period = 68 min

decay constant is given as

\lambda = 0.693/T = 0.693/68 = 0.0102

t = time period for decay = 2 hour 45 min = 2 (60) + 45 = 165 min

Ao = activity initially = ?

A = final activity = 15 mCi

using the equation

A = Ao e-\lambdat

inserting the values

15 = Ao e-((0.0102) (165))

Ao = 80.73 mCi